Let the the encryption function be f(a)=(3a+9)mod 26 . Encrypt "STOP THIEF" using this function.?

Answer is NQBE QGJXA. But i cant understand how this answer comes. I cant understand the process of solving this problem and get this correct answer.Please help me and solve this problem in detail.
1
Expert's answer
2012-10-16T09:40:59-0400
Assign numeric value to each letter of the alphabet:

A = 1 . . . . . H = 8 . . . . . . O = 15 . . . . . U = 21
B = 2 . . . . . I = 9 . . . . . . . P = 16 . . . . . V = 22
C = 3 . . . . . J = 10 . . . . . Q = 17 . . . . . W = 23
D = 4 . . . . . K = 11 . . . . . R = 18 . . . . . X = 24
E = 5 . . . . . L = 12 . . . . . S = 19 . . . . . Y = 25
F = 6 . . . . . M = 13 . . . . . T = 20 . . . . . Z = 26
G = 7 . . . . . N = 14

So in STOP THIEF, we get for each letter:

S = 19, f(19) = (3*19+9) mod 26 = 66 mod 26 = 14 = N
T= 20, f(20) = (3*20+9) mod 26 = 69 mod 26 = 17 = Q
O = 15, f(15) = (3*15+9) mod 26 = 54 mod 26 = 2 = B
P = 16, f(16) = (3*16+9) mod 26 = 57 mod 26 = 5 = E

T= 20, f(20) = (3*20+9) mod 26 = 69 mod 26 = 17 = Q
H = 8, f(8) = (3*8+9) mod 26 = 33 mod 26 = 7 = G
I = 9, f(9) = (3*9+9) mod 26 = 36 mod 26 = 10 = J
E = 5, f(5) = (3*5+9) mod 26 = 24 mod 26 = 24 = X
F = 6, f(6) = (3*6+9) mod 26 = 27 mod 26 = 1 = A

So "STOP THIEF" become "NQBE QGJXA"